OpenStudy (anonymous):

Solve for 3 ^2x = 27

OpenStudy (staceyg):

x=3

OpenStudy (staceyg):

do you want an explanation?

OpenStudy (anonymous):

Yes, please

OpenStudy (zzr0ck3r):

x is not 3 that would imply 3^6 = 26

OpenStudy (staceyg):

do you know what 3 to the second power is?

OpenStudy (anonymous):

It is 9.

OpenStudy (staceyg):

x=4.5?

OpenStudy (staceyg):

@zzr0ck3r

OpenStudy (anonymous):

It's not 4.5.. It said it was wrong

OpenStudy (zzr0ck3r):

\(3^{2x}=27\implies \log_3(3^{2x})=\log_3(27)\implies2x\log_3(3)=3\\\implies2x=3\implies x =\frac{3}{2}\)

OpenStudy (staceyg):

@zzr0ck3r So you're saying x= 1.5?

OpenStudy (staceyg):

@Soniahegner I thought it was 3^2x not 3^ 2x sorry

OpenStudy (anonymous):

It's okay, sorry for confusing you.

OpenStudy (zzr0ck3r):

oh i c.

OpenStudy (zzr0ck3r):

get used to writing things in fraction form, not decimal form. unless you are asked for the decimal.

OpenStudy (staceyg):

wait can you explain this to me? lol How would that equal 27? @zzr0ck3r

OpenStudy (zzr0ck3r):

if \(x=\frac{3}{2 }\) then \(3^{2x}=3^{2\frac{3}{2}}=3^3=\large27\)

OpenStudy (staceyg):

Oh that makes sense. When I was multiplying 2 by 3/2 and got 3, I multiplied 3 by 3 by accident. Thanks

OpenStudy (staceyg):

for explaining